# On Bessel’s Correction: Unbiased Sample Variance, the “Bariance,” and a Novel Runtime-Optimized Estimator

### v2

@@ Abstract
Bessel’s correction adjusts the denominator in the sample variance formula from $n$ to $n-1$ to produce an unbiased estimator for the population variance. This paper includes rigorous derivations, geometric interpretations, and visualizations. It then introduces the concept of “bariance,” an alternative pairwise distances intuition of sample dispersion without an arithmetic mean. Finally, we address practical concerns raised in Rosenthal’s article [1] advocating the use of $n$-based estimates from a more holistic $MSE$-based viewpoint for pedagogical reasons and in certain practical contexts. Finally, the empirical part using simulation reveals that the run-time of estimating population variance can be shortened when using an algebraically optimized “bariance“ approach to estimate an unbiased variance.
@@ 1 Introduction and Motivation
Variance estimation is a foundational task in statistics and econometrics, with the sample variance being the default estimator in most applications. The unbiased version, corrected by Bessel’s factor (dividing by $n-1$ rather than $n$), compensates for the loss of one degree of freedom due to pre-estimating the population mean. This correction is not just a simple algebraic trick–it admits deep geometric interpretations via orthogonal projections in $\mathbb{R}^{n}$ and can be derived rigorously from them.
@@
Despite its theoretical appeal, the unbiased estimator is not always the most optimal in practice. In small samples especially, its higher variance may lead to suboptimal inference. This has led researchers to consider shrunken estimators that intentionally trade off a small amount of bias for a significant reduction in variance, thereby minimizing mean squared error (MSE). For example, empirical Bayes methods shrink sample variances toward a global prior, stabilizing estimation across thousands of features in genomic studies [2]. Similar techniques based on James-Stein shrinkage have been explored for variance estimation in high-dimensional settings [3].
@@
Beyond the univariate case, shrinkage ideas are especially powerful in multivariate settings. In particular, shrinkage estimators for covariance matrices–such as the Ledoit-Wolf estimator [4]—have gained popularity in fields like econometrics and finance, particularly in the field of asset pricing. These estimators enhance the stability of sample covariance matrices by shrinking them toward structured targets (e.g., the identity matrix), significantly improving conditioning in high-dimensional models, which are known to break down [4]. This has practical relevance in the construction of variance-covariance matrices for portfolio optimization, factor models, and robust standard error estimation in large-scale regression analysis for econometric applications [5].
@@
In this broader context, this paper revisits classical variance estimation and introduces a novel perspective through the concept of an alternative measure of sample dispersion based on the average squared differences between all unordered pairs in a sample. It can be shown that for mean-centered data, the "bariance" equals exactly twice the unbiased sample variance. Moreover, a linear-time optimized formulation of the "bariance" can be derived using simple algebraic properties that avoids quadratic pairwise computation, making it both theoretically elegant and computationally efficient.
@@
Through a simulation study, I demonstrate that this optimized unbiased sample variance estimator remains unbiased and improves runtime. We then revisit the controversial idea—advocated by Rosenthal [1]—that dividing by $n$ (rather than $n-1$) may yield lower-MSE variance estimators in practice, especially when unbiasedness is not strictly required.
@@
This paper tries to bridge classical econometric and statistical theory with modern considerations of efficiency, robustness, and computational scalability, while re-instating the often under-estimated choices in estimator design or usage.
@@ 3 Derivation of Bias and Bessel’s Correction
We start by expanding the squared deviations:
@@
This shows that the estimator $S^{2}$ is biased, underestimating the population variance $\sigma^{2}$, because the denominator is larger then the numerator.
@@ 4 Geometric Interpretation of Estimated Variance and $n-1$ Degrees of Freedom
4.1 Orthogonal Projection
Let $\vec{X}\in\mathbb{R}^{n}$ be a data vector. Define the mean:
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$
and the mean vector:
$\vec{\mu}=\bar{X}\cdot\vec{1}$
We decompose:
$\vec{X}=\vec{\mu}+\vec{r},\quad\vec{r}=\vec{X}-\vec{\mu},\quad\vec{r}\in\vec{1}^{\perp}$
@@ 4.1 Orthogonal Projection
Figure 1: Projection of $\vec{X}$ onto mean direction $\operatorname{span}(\vec{1})$ and residual in $\vec{1}^{\perp}$
4.2 Dimension Argument
@@
$\vec{\mu}$ $=\bar{X}\cdot\vec{1}\in\operatorname{span}(\vec{1}),\quad\dim=1$
$\vec{r}$ $=\vec{X}-\vec{\mu}\in\vec{1}^{\perp},\quad\dim=n-1$
@@ 4.3 Sample Variance
The normalization by $n-1$ accounts for the loss of one degree of freedom due to estimation of the sample mean.
5 Introducing the “Bariance”
@@
This can be interpreted as the average squared length of all edges in the complete graph on the sample points.
@@
$\sum_{i\neq j}X_{i}X_{j}=\left(\sum_{i=1}^{n}\sum_{j=1}^{n}X_{i}X_{j}\right)-\sum_{i=1}^{n}X_{i}^{2}=\left(\sum_{i=1}^{n}X_{i}\right)^{2}-\sum_{i=1}^{n}X_{i}^{2}$
@@
$=2(n-1)\sum X_{i}^{2}-2\left(\left(\sum X_{i}\right)^{2}-\sum X_{i}^{2}\right)=2(n-1)\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}+2\sum X_{i}^{2}$
$=2n\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}$
@@
$\text{Bariance}=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}\equiv\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}$
See empirical verification in appendix A.
$\boxed{\text{Bariance}=\frac{2}{n-1}\left(\sum X_{i}^{2}\right)-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}}$
5.1 In the Case Of Mean-centered data
@@
$\text{Bariance}=\frac{2n}{n(n-1)}\sum X_{i}^{2}=\frac{2}{n-1}\sum X_{i}^{2}$
We now relate this to the unbiased sample variance:
$\hat{S}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}=\frac{1}{n-1}\sum X_{i}^{2}\quad\text{(since }\bar{X}=0\text{)}$
Therefore, following equality holds for the defined “Bariance“:
$\boxed{\text{Bariance}=2\cdot\hat{S}^{2}}$
This result shows that bariance represents twice the unbiased sample variance when the sample is mean-centered. It provides an elegant pairwise perspective on variance: instead of summing squared deviations from a central value, we sum squared differences between all pairs and average, despite the one we are currently looking from.
5.2 A Short Numerical Example with Five Numbers
Let $X=\{2,4,6,8,10\}\Rightarrow\bar{X}=6$
$\sum(X_{i}-\bar{X})^{2}=40,\quad\hat{S}^{2}=\frac{40}{4}=10,\quad\text{Bariance}=\frac{2\cdot 200}{20}=20$
Metric Value
Sample Mean $\bar{X}$ 6
Variance (biased) 8
Variance (unbiased) 10
Bariance 20
Bariance / 2 10
5.3 Graph-Theoretic View of Bariance
@@
Figure 2: Complete graph of sample values — each edge contributes as a component to the “bariance”.
5.4 Deviation from Mean (Variance) vs. Pairwise Differences (Bariance)
@@ 5.5 The Pairwise Difference Grid
Figure 4: Symmetric Grid of $(X_{i}-X_{j})^{2}$, $0.0$ for $i=j$, for $X=\{2,4,6,8,10\}$

### v5

# On Bessel’s Correction: Unbiased Sample Variance, the Bariance, and a Novel Runtime-Optimized Estimator

@@ Abstract
Bessel’s correction adjusts the denominator in the sample variance formula from $n$ to $n-1$ to ensure an unbiased estimator of the population variance. This paper provides rigorous algebraic derivations, geometric interpretations, and visualizations to reinforce the necessity of this correction. It further introduces the concept of Bariance, an alternative dispersion measure based on pairwise squared differences that avoids reliance on the arithmetic mean. Building on this, we address practical concerns raised in Rosenthal’s article [8], which advocates for $n$-based estimates from a mean squared error (MSE) perspective—particularly in pedagogical contexts and specific applied settings. Finally, the empirical component of this work, based on simulation studies, demonstrates that estimating the population variance via an algebraically optimized Bariance approach can yield a computational advantage. Specifically, the unbiased Bariance estimator can be computed in linear time, resulting in shorter run-times while preserving statistical validity.
@@
“This could be the end of it all,
released from the pain.
So self-indulgent and so insincere,
I’ll never bow to your lies again.”
— Megadeth, *Tornado of Souls*
The following notation is used throughout the paper:
$X_{1},X_{2},\dots,X_{n}\in\mathbb{R}$
Independent and identically distributed sample from a population.
$\mu=\operatorname{\mathbb{E}}[X_{i}]$ Population mean.
$\sigma^{2}=\operatorname{\mathrm{Var}}(X_{i})$ Population variance.
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ Sample mean.
$S^{2}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$ Biased sample variance estimator (denominator $n$).
$\hat{S}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$ Unbiased sample variance estimator (Bessel-corrected).
$\text{{Bariance}}=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}$ Pairwise variance estimator using all ordered pairs (each unordered pair counted twice).
$\text{{Bariance}}_{\text{opt}}=\frac{2n}{n(n-1)}\sum_{i=1}^{n}X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum_{i=1}^{n}X_{i}\right)^{2}$ Optimized scalar formula for Bariance.
$\sum_{i<j}(X_{i}-X_{j})^{2}$ Sum over all unordered pairwise squared differences (each pair counted once).
Note: $\sum_{i\neq j}(X_{i}-X_{j})^{2}=2\sum_{i<j}(X_{i}-X_{j})^{2}$.
$\text{Bias}(\hat{\theta})=\operatorname{\mathbb{E}}[\hat{\theta}]-\theta$ Bias of an estimator $\hat{\theta}$.
$\text{Var}(\hat{\theta})=\operatorname{\mathbb{E}}[(\hat{\theta}-\operatorname{\mathbb{E}}[\hat{\theta}])^{2}]$ Variance of an estimator.
$\text{MSE}(\hat{\theta})=\text{Var}(\hat{\theta})+\text{Bias}^{2}(\hat{\theta})$ Mean squared error of an estimator.
$\vec{X}\in\mathbb{R}^{n}$ Sample vector in Euclidean space.
$\vec{1}\in\mathbb{R}^{n}$ All-ones vector.
$\vec{r}=\vec{X}-\vec{\mu}$ Residual vector after projection onto the mean.
$\vec{r}\in\vec{1}^{\perp}$ Residual lies in the orthogonal complement of the mean vector.
$\Gamma(k,\theta)$ Gamma distribution with shape $k$ and scale $\theta$.
$\mathcal{N}(\mu,\sigma^{2})$ Normal distribution with mean $\mu$ and variance $\sigma^{2}$.
$\sum X_{i}$ Sum of sample values.
$\sum X_{i}^{2}$ Sum of squared sample values.
Note: The term Bariance emphasizes variance computed from pairwise squared differences instead of deviations from the mean.
@@ 1 Introduction and Motivation
Variance estimation is a foundational task in statistics and econometrics, with the sample variance being the default estimator in most applications. The unbiased version, corrected by Bessel’s factor (dividing by $n-1$ rather than $n$), compensates for the loss of one degree of freedom due to pre-estimating the population mean. This correction is not just a simple algebraic trick—it admits deep geometric interpretations via orthogonal projections in $\mathbb{R}^{n}$ and can be derived rigorously from them.
@@
Despite its theoretical appeal, the unbiased estimator is not always the most optimal in practice. In small samples especially, its higher variance may lead to suboptimal inference. This has led researchers to consider shrunken estimators that intentionally trade off a small amount of bias for a significant reduction in variance, thereby minimizing mean squared error (MSE). For example, empirical Bayes methods shrink sample variances toward a global prior, stabilizing estimation across thousands of features in genomic studies [9]. Similar techniques based on James–Stein shrinkage have been explored for variance estimation in high-dimensional settings [4].
@@
Beyond the univariate case, shrinkage ideas are especially powerful in multivariate settings. In particular, shrinkage estimators for covariance matrices—such as the Ledoit–Wolf estimator [6]—have gained popularity in fields like econometrics and finance, particularly in the field of asset pricing. These estimators enhance the stability of sample covariance matrices by shrinking them toward structured targets (e.g., the identity matrix), significantly improving conditioning in high-dimensional models, which are known to perform poorly [5]. This has practical relevance in the construction of variance-covariance matrices for portfolio optimization, factor models, and robust standard error estimation in large-scale regression analysis for econometric applications.
@@
In this broader context, this paper revisits classical variance estimation and introduces a novel perspective via an alternative measure of sample dispersion based on the average squared differences between all unordered pairs in a sample. We formally define this estimator as the Bariance, a term that reflects its construction from pairwise distances rather than deviations from a mean. It can be shown that for mean-centered data, the Bariance equals exactly twice the unbiased sample variance. Moreover, a linear-time optimized formulation of the Bariance can be derived using simple algebraic properties that avoids quadratic pairwise computation, making it both theoretically elegant and computationally efficient.
Although the pairwise difference approach has roots in classical statistics—such as U-statistics [7], dissimilarity-based dispersion measures, and even the Gini coefficient [2]—the contribution here is a novel, unbiased estimator that is computationally optimized for runtime efficiency. In this respect, Bariance bridges theoretical variance estimation with algorithmic efficiency, a consideration critical in big data contexts, real-time systems, and streaming analytics.
While computational efficiency is one of its key advantages, the Bariance measure may also prove valuable in applied scenarios where the concept of central tendency is unstable, ill-defined, or misleading. For example, in domains such as network analysis, genomics, ordinal survey research, or clustering, statistical dispersion is often better captured through relational or pairwise structures rather than deviations from a single global mean. In such contexts, the Bariance shares conceptual kinship with the Gini coefficient, which also operates on pairwise differences but in a distributional inequality framework. Unlike Gini, however, Bariance preserves unbiasedness for variance estimation under i.i.d. sampling and scales naturally in high-dimensional or streaming environments. These features make it particularly attractive for modern applications in unsupervised learning, robust statistics, and high-throughput data pipelines—where traditional variance measures may either fail or become computationally prohibitive.
@@
Through an empirical simulation study, I demonstrate that this optimized unbiased sample variance estimator remains unbiased and improves runtime. The simulated empirical runtimes section includes confidence intervals, hardware specifications, seed initialization, and multiple replications, thereby addressing robustness, reproducibility, and statistical reliability. Furthermore, Appendix C replicates the result in a local hardware environment using an alternative high-level programming language. We then revisit the controversial idea—advocated by Rosenthal [8]—that dividing by $n$ (rather than $n-1$) may yield lower-MSE variance estimators in practice, especially when unbiasedness is not strictly required.
@@
To sum up, the Bariance framework bridges computational efficiency with applied relevance, offering a theoretically grounded yet practically flexible alternative to traditional variance estimators. This paper thus aims to bridge classical econometric and statistical theory with modern considerations of efficiency, robustness, and computational scalability, while highlighting the often underestimated choices in estimator design or usage.
@@ 3 Derivation of Bias and Bessel’s Correction
Expand the squared deviations:
@@
This shows that the estimator $S^{2}$ is biased, underestimating the population variance $\sigma^{2}$, because the denominator is larger than the numerator.
@@ 4 Geometric Interpretation of Estimated Variance and $n-1$ Degrees of Freedom
Orthogonal Decomposition
Let $\vec{X}:=\begin{bmatrix}X_{1}&\dots&X_{n}\end{bmatrix}^{\top}\in\mathbb{R}^{n}$. Define the mean
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i},\quad\vec{\mu}=\bar{X}\cdot\vec{1}.$
Then
$\vec{X}=\vec{\mu}+\vec{r},\quad\vec{r}:=\vec{X}-\vec{\mu},\quad\vec{r}\in\vec{1}^{\perp}.$
Indeed, $\vec{1}^{\top}\vec{r}=\sum_{i=1}^{n}(X_{i}-\bar{X})=0$.
@@
Figure 1: Orthogonal decomposition of $\vec{X}$
Dimension and Degrees of Freedom
@@
$\vec{\mu}$ $\in\operatorname{span}(\vec{1}),\quad\dim=1$
$\vec{r}$ $\in\vec{1}^{\perp},\quad\dim=n-1$
@@
Common Application using Orthogonal Decomposition for Dimensionality reduction: Principal Component Analysis (PCA)
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ be a data matrix with rows as observations and columns as variables. Let $\tilde{X}:=X-\vec{1}\bar{x}^{\top}$, where $\bar{x}=\frac{1}{n}X^{\top}\vec{1}\in\mathbb{R}^{p}$.
PCA seeks orthonormal vectors $\vec{v}_{i}\in\mathbb{R}^{p}$ satisfying:
$\tilde{X}^{\top}\tilde{X}\vec{v}_{i}=\lambda_{i}\vec{v}_{i},\quad i=1,\dots,p.$
Then $\vec{z}_{i}:=\tilde{X}\vec{v}_{i}$ are commonly known as $i$th principal components ($PC_{i}$), with variances $\operatorname{Var}(\vec{z}_{i})=\lambda_{i}/(n-1)$.
5 Introducing the Bariance and an optimized linear Estimator
@@
The term Bariance is simply chosen to emphasize the estimator’s foundation on pairwise between-sample variance rather than deviations from the mean, highlighting its construction from pairwise squared differences. This can also be interpreted as the average squared length of all edges in a complete graph on the sample points.
@@
$\sum_{i\neq j}X_{i}X_{j}=\left(\sum_{i=1}^{n}\sum_{j=1}^{n}X_{i}X_{j}\right)-\sum_{i=1}^{n}X_{i}^{2}=\left(\sum X_{i}\right)^{2}-\sum X_{i}^{2}$
@@
$=2(n-1)\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}+2\sum X_{i}^{2}=2n\sum X_{i}^{2}-2\left(\sum X_{i}\right)^{2}$
@@
$\text{{Bariance}}=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}\equiv\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.$
For empirical verification of this algebraic identity see Appendix A.
$\boxed{\text{{Bariance}}_{\text{opt}}:=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.}$
5.0.1 In the Case Of Mean-centered data
@@
$\text{{Bariance}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}=\frac{2}{n-1}\sum X_{i}^{2}.$
We now relate this to the unbiased sample variance estimator:
$\hat{S}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}=\frac{1}{n-1}\sum X_{i}^{2}\quad\text{(since }\bar{X}=0\text{)}.$
Therefore, the following equality holds for the defined “Bariance“:
$\boxed{\text{{Bariance}}=2\cdot\hat{S}^{2}.}$
This result shows that Bariance represents twice the unbiased sample variance when the sample is mean-centered. It provides an elegant pairwise perspective on variance: instead of summing squared deviations from a central value, we sum squared differences between all pairs and average, regardless of the reference point within the sample.
5.1 Properties of the Bariance and Estimator Comparison
Let $\theta:=\sigma^{2}=1$. Then:
$\operatorname{\mathbb{E}}[\hat{S}^{2}]$ $=\theta,$
$\operatorname{\mathbb{E}}[\text{{Bariance}}]$ $=2\theta,$
$\text{Bias}(\hat{S}^{2})$ $=0,$
@@ 5.1 Properties of the Bariance and Estimator Comparison
$\operatorname{\mathrm{Var}}(\hat{S}^{2})$ $=\frac{2\theta^{2}}{n-1},$
$\operatorname{\mathrm{Var}}(\text{{Bariance}})$ $=4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2}),$
$\text{MSE}(\hat{S}^{2})$ $=\operatorname{\mathrm{Var}}(\hat{S}^{2}),$
[…]
Subsequent numerical verification of theoretical relationships with $n=100$ and $\tau=1000$
Numerically (for $n=100$, $\mathcal{N}(0,1)$, $\tau=1000$):
Estimator Point Estimate Bias Variance MSE
Unbiased sample variance estimator 1.00091 0.00091 0.02156 0.02151
Bariance 2.00181 1.00181 0.08625 1.08968
$\operatorname{\mathrm{Var}}(\hat{S}^{2})=0.01988,\quad\operatorname{\mathrm{Var}}(2\cdot\hat{S}^{2})=0.07951,$
$\text{MSE}(\hat{S}^{2})=0.02151,\quad\text{MSE}(2\cdot\hat{S}^{2})=1.0832,$
$4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})=0.07951\quad\Rightarrow\quad\text{Var identity approximately holds},$
$4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})+\theta^{2}=0.07951+1=1.07951\quad\Rightarrow\quad\text{MSE identity approximately holds}.$
[…]
Summary:
  • $\hat{S}^{2}$: unbiased, lower MSE, standard estimator.
  • Bariance: biased (by $+\theta$), higher MSE due to both variance inflation and squared bias.
  • Useful relation: $\text{{Bariance}}=2\cdot\hat{S}^{2}$. (Because: $\theta$ = ($\theta$ + $\theta$)/2)
5.2 A Graph-Theoretic View of Bariance
[…]
Figure 2: Complete graph over sample points. Each edge represents a pairwise squared difference that contributes to the Bariance.
Each dashed edge on the graph corresponds to a pair $(X_{i},X_{j})$, and the squared difference $(X_{i}-X_{j})^{2}$ can be viewed as an edge weight. Since the graph is fully connected (a complete graph $G_{n}$), there are $\binom{n}{2}$ such edges.
From a graph-theoretic standpoint, Bariance is the average squared edge length of the complete weighted graph over the sample. In this context:
- Nodes = observations $X_{i}$
- Edges = pairwise differences
- Weights = squared differences $(X_{i}-X_{j})^{2}$
This perspective connects naturally to distance-based dispersion measures in statistical graph theory, including energy statistics and certain U-statistics. It also provides an intuitive, coordinate-free alternative to the variance’s reliance on a central location (i.e., the mean).
5.3 Deviation from Mean (Variance) vs. Pairwise Differences (Bariance Components)
@@ 5.4 The Pairwise Difference Grid
Figure 4: Symmetric matrix of squared pairwise differences $(X_{i}-X_{j})^{2}$ for $X=\{2,4,6,8,10\}$. Diagonal elements are zero (self-distances).
This matrix visualizes all pairwise squared differences between elements of $X$. Its key properties:
- Symmetry: $(X_{i}-X_{j})^{2}=(X_{j}-X_{i})^{2}$, so the grid is symmetric across the main diagonal.
- Diagonal: Always zero, since $(X_{i}-X_{i})^{2}=0$.
- Off-diagonal structure: Each non-zero entry contributes to the total in the Bariance calculation.
This grid-based view reinforces how the redundancy in the pairwise differences allows us to simplify the computation algebraically. Instead of evaluating all $n(n-1)$ pairs, we can summarize the matrix using aggregate row/column sums, which enables the linear-time algebraically optimized derived formula for the Bariance:
$\text{{Bariance}}_{\text{opt}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}.$
Thus, symmetry is not just a visual feature—it underpins the algebraic transformation that reduces the quadratic computational complexity to linear.
5.5 Numerical Verification of Bariance-Estimator properties using Gamma-distributed Data
Let $X\sim\Gamma(k=2,\theta=2)\Rightarrow\operatorname{\mathbb{E}}[X]=4,\quad\operatorname{\mathrm{Var}}(X)=8$, Using $n=100$, $\tau=1000$ we obtain:
Summary Table of Empirical Bariance Point Estimator Performance
Estimator Mean Bias Variance MSE
Unbiased Sample Variance Estimator 8.00087 0.00087 2.92574 2.92281
Bariance 16.00174 8.00174 11.70295 75.71907
Numerical Verification of Theoretical Relationships
Let $\sigma^{2}:=8\Rightarrow\sigma^{4}=64$:
$\operatorname{\mathrm{Var}}(\hat{S}^{2})=2.92574,\quad\operatorname{\mathrm{Var}}(\text{{Bariance}})=11.70295,$
[…]
We numerically verify:
$\operatorname{\mathrm{Var}}(\text{{Bariance}})\approx 4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})\quad\Rightarrow\quad 11.70295\approx 4\cdot 2.92574=11.70296,$
$\text{MSE}(\text{{Bariance}})\approx 4\cdot\operatorname{\mathrm{Var}}(\hat{S}^{2})+\sigma^{4}\quad\Rightarrow\quad 75.71907\approx 4\cdot 2.92574+64=75.70295,$
[…]